OpenStudy (anonymous):

A catcher stops a ball travelling at 26.2 m/s in a distance of 85.1 cm and feels a force of 313 N against his glove. What is the mass of the ball?

OpenStudy (stormfire1):

first find the acceleration: \[Vf^2 - Vi^2 = 2a(Xf-X i)\]\[(0m/s)^2 - (26.2 m/s)^2 = 2a(0 - 0.851m)\]\[\frac{-686.44}{2} = -343.22 = a(-0.851m)\]\[\frac{-343.22}{-0.851m}=~403 m/s^2\] Since force = mass times acceleration: \[313N = mass * 403m/s^2\]\[\frac{313N}{403} m/s^2 = 0.777kg\]

OpenStudy (stormfire1):

In the last statement, the m/s^2 on the left should be in the denominator of the fraction...I guess I made a typo